OpenStudy (anonymous):

How many grams of chlorine gas (Cl2) are in a 17.8 liter sample at 1.1 atmospheres and 29 degrees C? can someone help me understand this...im not sure how to do it, i think it might be the percent catching me off guard...can someone explain to me how to get the answer?

OpenStudy (anonymous):

@Rushwr can you help me with this?

OpenStudy (mayankdevnani):

use PV=nRT

OpenStudy (rushwr):

let me tell u the easy way to solve these first write down the data u are given with We know the volume (V), Pressure (P) and temperature (T) R is the universal gas constant now u can see we know all the data needed except no. of moles (n) in PV=nRT equation right? We know moles = mass divided by molar mass Molar mass of Cl2 is approximately 71gmol^-1

OpenStudy (rushwr):

first find the moles or else u can directly substitute m/M instead of n in the equation m= mass M= molar mass

OpenStudy (rushwr):

Can u try and get the answer here !

OpenStudy (anonymous):

ok give me a sec and ill try :)

OpenStudy (anonymous):

how do i get the mass?

OpenStudy (rushwr):

moles = mass divded by molar mass moles is the n in the equation instead of PV=nRT we can write the equation like PV=(m/M)RT right?

OpenStudy (rushwr):

P = pressure = 1.1 atm V = volume = 17.8 L n = number of moles Cl2 (unknown) R = gas constant = 0.08206 L*atm/mol*K T = temperature = 29C + 273.15 = 302.15 K

OpenStudy (anonymous):

ok that kinda makes sense...

OpenStudy (rushwr):

these are the data. Make sure that u convert ur temperature to kelvin

OpenStudy (anonymous):

where did the 273.15 come from?

OpenStudy (rushwr):

\[PV= nRT\] \[n= \frac{ m }{ M }\] \[PV= \frac{ m }{ M }RT\]

OpenStudy (rushwr):

When u convert a temperature from Celsius to kelvin u need to add 273.15 to the Celsius value to convert t to kelvin

OpenStudy (rushwr):

I hope that makes sense. oh u can also add 273 approximately

OpenStudy (anonymous):

so i would use that number every time i converted celcius to kelvin or just for this?

OpenStudy (rushwr):

yes exactly

OpenStudy (rushwr):

@AlexisIB Make sure that u are thorough with the conversions

OpenStudy (anonymous):

ok what do i do with everything now? do i multiply?

OpenStudy (anonymous):

do i keep it like it is?

OpenStudy (rushwr):

now substitute the values

OpenStudy (anonymous):

ok...not to sound stupid but how do i do that? :(

OpenStudy (rushwr):

PV= nRT right? First we are gonna find moles (n) We know pressure (P) = 1.1 atm Volume (V) = 17.8L Universal gas constant (R) = 0.08206 L*atm/mol*K Temperature (T) = 29+273 = 302K

OpenStudy (rushwr):

now get the value for n !

OpenStudy (anonymous):

ok but i still dont know how...but ill call my teacher, thanks for helping me though

OpenStudy (rushwr):

no problem . And sorry for late replies my OS is lagging

OpenStudy (anonymous):

@Rushwr i think i figured it out without calling my teacher...can you check to see if im right so far?

OpenStudy (anonymous):

@mayankdevnani can you check?

OpenStudy (anonymous):

P=1.1 V=17.8 R=.0821 T=29' C + 273.15 = 302.15 K (always use 273.15 when converting from degrees to kelvin) 1.1 x 17.8 / 0.0821 x 302 =0.7895 Mass= .7895 mol x 70.906 g/mol

OpenStudy (mayankdevnani):

yup ! good job :) @AlexisIB

OpenStudy (rushwr):

I'm sorry I was a little late for this. Anyways yeah great job @AlexisIB

OpenStudy (anonymous):

thank you!!! im so happy that i actually did it!!

OpenStudy (rushwr):

:)

OpenStudy (anonymous):

both of you are awesome :)

OpenStudy (rushwr):

lol thanks

OpenStudy (mayankdevnani):

you are awesome @AlexisIB

OpenStudy (mayankdevnani):

YOUR HARDWORK PAYS OFF !

OpenStudy (anonymous):

Thanks!!

OpenStudy (mayankdevnani):

your welcome :)